# Boosting Performance on ARC is a Matter of Perspective

### v1

@@ 6 Discussion
[…]
By applying these ideas to ARC, we underline a broader principle: when dealing with structured or abstract reasoning tasks, the key factor is to exploit valid semantic-preserving transformations, forcing a model to remain consistent across multiple views of the same problem. This allows us to use a single model as an ensemble of experts. We believe this perspective can generalize [^2] The appendix provides preliminary experiments that test our setup with largely unchanged hyperparameters, solving Sudoku puzzles with 50% missing values at a success rate of 97.5% compared to a greedy baseline solution of 70.5%. to more complex symbolic reasoning challenges, wherever such transformations can be defined. Our results demonstrate that large language models, properly steered in inference and supported by prior aware scoring, can go beyond default sampling approaches to capture deeper structures in abstract domains.
Acknowledgement
We would like to express our sincere gratitude to Lambda, for providing computational resources essential for optimizing our preliminary training phase within our overall training pipeline. Specifically, they supplied us with a server equipped with 8xH100 GPUs, enabling rapid iteration on our ideas. Their support was instrumental for reaching our final score.
This work has been supported by the ”Research Center for Algorithmic Intelligence as an Emergent Phenomenon” (funded by the Carl-Zeiss-Stiftung) and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), project 233630050 (Collaborative Research Center TRR 146).
Generative AI language tools were used for text editing in this manuscript; all AI-generated output was subsequently reviewed, revised, and validated by the authors, who assume full responsibility for the final manuscript.
@@ Acknowledgement
Appendix A Proofs
Theorem A.1 (Error Bound for Log-Pooled Augmentations).
@@ Appendix A Proofs
A.1 Sudoku Experiment (Preliminary)
We create Sudokus by first removing 50% of the cells from a valid puzzle, rejecting any puzzle that does not have a unique solution. In total, 2000 such puzzles serve as training data, while 400 are used for testing. We structure these 400 puzzles into tasks that each contain five training example pairs and one test puzzle, mirroring the ARC task format and enabling the reuse of the same hyperparameters from the main text. Since the entire dataset is available at train-time, we do not apply Test-Time-Training (TTT) and instead train for only 32 epochs. We employ the same augmentation procedures described in the main text (i.e., all dihedral group transformations and color shuffling), yielding 16 unique augmentations per puzzle. Although these results are preliminary, they demonstrate the consistency of our approach when reusing the same hyperparameters: the baseline strategy of selecting the most probable solution achieves 70.5% accuracy, while our product of experts approach achieves 97.5%.

### v2

@@ 5.4 Selection Strategies
5.5 ConceptARC
To make sure we do not overfit on the original ARC data, we further evaluate our method on ConceptARC (Moskvichev et al., 2023) - an ARC-like dataset containing tasks sorted into specific conceptual categories. Our method achieves 73.3% 2-guess accuracy on ConceptARC (using the exact same hyperparameters as DFS T=9%), showing that we generalize well to other ARC-like datasets of similar difficulty.
Figure 6: Results of the Sudoku experiments (plot equivalent to Figure 4, but showing top-1 accuracy instead of top-2). We can see that our product of experts approach increases the accuracy substantially to 53% solved Sudoku puzzles over simply selecting the generated solution with the highest sampling probability. Note that generation accuracy completely coincides with product of experts probability, showing that if a correct solution is sampled, our approach consistently selects it.
5.6 Sudoku
We further test our approach on the Sudoku 3M dataset (Radcliffe, 2020) to evaluate generalizability of the method to different domains. Since the underlying ”rules” of Sudoku remain consistent between tasks, we do not use any test-time training in this case. Instead, we start out with our Llama 3B model pre-trained on ARC, which we then finetune again on $128000$ Sudoku tasks. As the Sudoku tasks never have any ambiguity, we report top-1 accuracy rather than top-2. To handle the increased complexity for the LLM compared to ARC, we use DFS with a threshold of T=1%, which provides a good trade-off between accuracy and runtime (see Figure 6). This setup reaches $53\%$ accuracy on $1000$ randomly chosen unseen Sudoku puzzles, far better than state-of-the-art LLMs, which have a solve-rate less than 3% on comparable benchmarks (Seely et al., 2025). Notably, if the correct solution of a puzzle is sampled, we select it in $100\%$ of cases. This is caused by the fact that Sudoku correctness is simple to evaluate. Using our standard augmentations described in Section 5.3 on the predictions, the model can identify errors more frequently, thereby significantly reducing the likelihood of false positives.
@@ 6 Discussion
Despite ARC’s reputation for complexity, our two-phase “generate-then-re-score” routine achieves SOTA results among open models. While only a single closed-source solution (arcprize.org, 2025) posts a higher absolute score at \$17 per task, our fully open-source process stands out for its transparency, reproducibility and, above all, its cost-effectiveness of only $0.02\$$ per task.
By applying these ideas to ARC, we underline a broader principle: when dealing with structured or abstract reasoning tasks, the key factor is to exploit valid semantic-preserving transformations, forcing a model to remain consistent across multiple views of the same problem. This allows us to use a single model as an ensemble of experts. We believe this perspective can generalize to more complex symbolic reasoning challenges, wherever such transformations can be defined. Our results demonstrate that large language models, properly steered in inference and supported by prior aware scoring, can go beyond default sampling approaches to capture deeper structures in abstract domains.
6.1 Future Work
Building upon our insights, several promising directions emerge for future investigation. First, it would be valuable to further explore the generalizability of using a single large language model as a Product-of-Experts through augmentations beyond ARC-specific transformations. In particular, text-based augmentations such as linguistic reformulations or stylistic variations present possible paths to extend our method to a broader array of natural language reasoning tasks. Second, the effectiveness of our depth-first search (DFS) candidate-generation strategy warrants evaluation beyond ARC-like puzzles; exploring tasks such as logical reasoning, program synthesis, or mathematical problem-solving could yield insights into its broader applicability and effectiveness in structured problem-solving domains.
@@ 6.1 Future Work
Appendix A Appendix: DFS Algorithm
The algorithm presented here assumes that the model supports internal caching for already seen sequences and only needs to process the newly added tokens. Note that we use negative log probabilities to avoid numerical issues, while the main paper uses percentage values for clarity.
Our actual implementation differs from this simple variant, as we are using unsloth, which does not support dynamic caching and requires us to prune the key-value-cache of the transformer ourselves.
Furthermore, we use various performance optimizations, like a simultaneous initial forward pass of the best known sequence including prompt and prediction (which is much faster than token-by-token generation) as well as aggregating the sequences during backtracking to avoid the unnecessary processing of sequences that would be discarded later
Algorithm 1 Depth-First Probability-Guided Sampling for LLMs.
[$model,prompt,threshold,max\_len,eos\_id$]DFS_sample
  Input: $model$ is the language model
  Input: $prompt$ is the prompt that should be completed
  Input: $threshold$ is the maximum negative log probability allowed
  Input: $max\_len$ is the maximum length (including the prompt)
[…]
  if $tokens[-1]=eos\_id$ or $|tokens|\geq max\_len$ then
     return $(score,tokens)$
  end if
  $next\_token\_logits\leftarrow model.predict\_logits(tokens)[-1]$
  $next\_token\_log\_prob\leftarrow-\text{log\_softmax}(next\_token\_logits)$
  $valid\_sequences\leftarrow\emptyset$
  for each possible next token $t$ do
     $next\_score\leftarrow score+next\_token\_log\_prob[t]$
     if $next\_score\leq threshold$ then
        $next\_tokens\leftarrow current\_tokens+[t]$
        $continuations\leftarrow\text{Explore}({next\_tokens,next\_score})$
        $valid\_sequences\leftarrow valid\_sequences\cup continuations$
     end if
  end for
  return $valid\_sequences$ENDFUNCTION
[…]
Appendix B Appendix: Training Parameters
Table 4: Training parameters and times for the initial and the test-time fine-tuning processes. Test-time fine-tuning is performed separately for each task, each time starting from the initially fine-tuned base model.
Initial Fine-Tuning Test-Time Fine-Tuning
Batch size 4 1
Gradient acc. steps 2 1
LoRA rank 256 32
LoRA $\alpha$ 24 16
LoRA bias off off
rank-stabilized LoRA on on
LR (LoRA adapters) $1\mathrm{e}{-4}$ $1\mathrm{e}{-4}$
LR (embeddings) $1\mathrm{e}{-5}$ $1\mathrm{e}{-5}$
LR schedule cosine cosine
LR warmup phase 25% 50%
Weight decay off off
Optimizer adamw_8bit adamw_8bit
Base model quantization 4 bit 4 bit
Data type bfloat16 bfloat16
Trained tokens outputs only outputs only
Training dataset RE-ARC single task examples
Number of Epochs 368 [Llama] 1200 [NeMo] 64
Training performed on 1x Nvidia H100 1x Nvidia RTX 4090
Training time 15 hrs. [Llama] 98 hrs. [NeMo] 12 sec./task [Llama] 51 sec./task [NeMo]
Table 5: Reduced Token Set for ARC-AGI-specific LLM Model
Token Category Tokens Purpose
Alphabet A-Z, a-z (excl. I,O,i,o) Learned pre-prompt tokens
Numbers 0-9 Encoding the 10 colors
[…]
Input/Output I, O Signals start of problem input/output
Begin token $\langle bos\rangle$ Inserted once at the beginning
End token $\langle eos\rangle$ Inserted after each output
Padding token $\langle pad\rangle$ Internal usage (e.g. batching)
Appendix C Appendix: Product of Experts Proof
Theorem C.1 (Error Bound for Log-Pooled Augmentations).